Tech Meeting Minutes 20121406
|Table of contents|
Present: Sam, Gareth, Andy, Mike, Dave (chair & minutes)
SE real problem. Thought this was sorted, seems to have fallen over again.
Hardware damage after power cut
power supplies in worker nodes
some disk servers
Temperature sensors to autoshutdown workers have been deployed and tested
There have now been meetings discussing adding the chiller to the backup generator
The power being used now is an alternate feed. There will be another downtime when the feed has to be switched back
Need A/C in room with UPS
Nothing to report, Andy's getting caught back up
Hot spots still in evidence, transition from SL4-SL5 almost complete which can be followed by file rebalancing.